OpenStudy (anonymous):

1. Write the equation for the vertical line passing through (5, 8). A. x = 5 B. x = 8 C. x = 0 D. x = 5/8 2. Write the equation for the horizontal line passing through (4, -6). A. y = 4 B. y = -6 C. y = 4/6 D. y = -6/4 3 . Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1) 4. Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

OpenStudy (vishal_kothari):

1.x=5

OpenStudy (vishal_kothari):

2.y=-6

OpenStudy (anonymous):

Really? Want whole homework problem list done for you?

OpenStudy (anonymous):

You're shameless

OpenStudy (anonymous):

Vishal - that is not helping anyone. May be explaining how to get their own answers is a better approach to try?

OpenStudy (anonymous):

Sorry I'm not quite sure that is the right word.

OpenStudy (anonymous):

Well then can u help me?

OpenStudy (anonymous):

First, please post each question separately.

OpenStudy (anonymous):

okay, wen i used to do tht evryone got confused and told me to post them all in one post. soo....

OpenStudy (anonymous):

Do you want to start with the first one?

OpenStudy (anonymous):

No we can start with 3.

OpenStudy (anonymous):

Well the point slope equation of a line looks like this:\[y = mx+b\]

OpenStudy (anonymous):

You know that: m = 6 and \[(x,y)=(5,-1)\]

OpenStudy (anonymous):

-1 = 6(5) + b

OpenStudy (anonymous):

Solve for b and you get. b = -31

OpenStudy (anonymous):

And that I all you need.

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

Try to focus on one problem at a time.

OpenStudy (anonymous):

A little.

OpenStudy (anonymous):

What part is confusing you?

OpenStudy (anonymous):

So will it be A?

OpenStudy (anonymous):

Not it will be D?

OpenStudy (anonymous):

yeah! You're right. Is D.

OpenStudy (anonymous):

Good work =)

OpenStudy (anonymous):

Thank you.Okay can u help me with 4?

OpenStudy (anonymous):

What is D? That is wrong.

OpenStudy (anonymous):

3 is D

OpenStudy (anonymous):

Yes, sorry is C, I confused the letters

OpenStudy (anonymous):

NO! It is C.

OpenStudy (anonymous):

is (y+1)=6(x-5)

OpenStudy (anonymous):

okay can u help me with 4 now?

OpenStudy (anonymous):

Do you see why?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thts not helping me thts telling me the answer. I wnt help.

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

You need to do operations in order to the equation look like one of the options.

OpenStudy (anonymous):

Okay well i kind of figured it out on my own. and now can u help me with question number 4 on the other one tht i posted?

OpenStudy (anonymous):

Okay